Variable Cost per Visit $10.00

Projected Number of Visits 10000

Total Variable Cost $100,000

Annual Direct Fixed Cost $500,000.00

Overhead $50,000.00

Total Cost $650,000

Desired Profit $650,000

What is the price per visit to reach a desired profit of $650.00?
Answer #1
Total fixed cost = Annual direct fixed cost + Overhead = 500000 + 50000 550000
Projected number of visits = ( Total fixed cost + Desired profit ) / Contribution margin per visit
10000 = ( 550000 + 650000 ) / Contribution margin per visit
10000 = 1200000 / Contribution margin per visit
Contribution margin per visit = 1200000 / 10000 120.00
Price per visit = Contribution margin per visit + Variable cost per visit = 120.00 + 10.00 130.00